In 700 c.c of a mixture of three liquids volume of 1st and 2nd are in a ratio 2:3 and the volumes of 2nd and 3rd are in a ratio 4:5. How much of 1st and 3rd be mixed with the mixture so as to reach the ratio to 5:6:10.

Answer by ankor@dixie-net.com(22740)

In 700 c.c of a mixture of three liquids volume of 1st and 2nd are in a ratio 2:3
and the volumes of 2nd and 3rd are in a ratio 4:5.
How much of 1st and 3rd be mixed with the mixture so as to reach the ratio to 5:6:10.
:
All we need to do is find amts that will give us the resulting mixture
:
Find the multiplier (x) for the resulting mixture which is: 5:6:10
5x + 6x + 10x = 700
21x = 700
x = 700/21
x = 33.33
then amt of the 3 liquids are:
5*33.33 = 166.67 cc of the 1st liquid
6*33.33 = 200.0 cc of the 2nd liquid
10*33.33 = 330.33 cc of the 3rd liquid
Adds up to 700 cc